OpenStudy (anonymous):

Diophantus' youth lasted 1/6 of his life. He grew a beard after 1/12 more of his life. After 1/7 more of his life, Diophantus married. Five years later he had a son. The son lived exactly 1/2 as long as his father, and Diophantus died just four years after his son's death. All of this totals the years Diophantus lived." NEED HELP..

OpenStudy (alekos):

by my calculation he lived 84 years. do you have an answer?

OpenStudy (anonymous):

How did you get 84?

OpenStudy (alekos):

Following the script... L = L/6 + L/12 + L/7 + 5 + L/2 + 4 Then i solved for L (no of years of Diophantus' life)

OpenStudy (alekos):

happy with that?

OpenStudy (anonymous):

I got 84 also. (1/6)x + (1/12)x + (1/7)x 5 + (1/2)x +4 = x Simplify: (3/28)x = 9 x = 84 years old

OpenStudy (alekos):

yep, looks good. thats what it is

OpenStudy (anonymous):

Thanks.

OpenStudy (alekos):

your welcome